# Scalable Hypergraph Structure Learning with Diverse Smoothness Priors

### v1

@@ III HGSI
calculated element-wise on $\mathbf{z}$ where $\mathbf{w}^{\star}\in(0,1]^{\hat{D}}$. The recovery of $\mathbf{w}^{\star}$ is the equivalent of learning the hypergraph structure since it contains the weights for each potential hyperedge, indicating how confident we are that each hyperedge is in the hypergraph. We note that the length of $\mathbf{w}^{\star}$ is not $D$, but rather $\hat{D}$. It is recognized that computing $\mathbf{w}^{\star}$ is time inefficient and, as $N$ and $M$ increase, often completely infeasible due to the vast number of hyperedge permutations. So, to limit the number of hyperedge possibilities, HGSI employs the use of a K-NN based method for possibility reduction. Specifically, given $\mathcal{K}$, determine the $k_{l}-1$ nearest neighbors of each node based on time-series signal observations. Then, let the combinations of the original node and its nearest neighbors be the new set of possible hyperedges where $\hat{D}$ is the number of hyperedge possibilities in the set reduced by K-NN.
@@
HGSI has advantages when it comes to scalability through the use of a closed form solution and the hyperedge reduction technique. However, both of these factors also cause certain shortcomings as they relate to assumptions on hypergraph structure. The first is that the format of (10) improperly constrains the set of learned hyperedges in a non-uniform hypergraph structure as a factor of $\mathbf{z}$. So, if there is some non-uniform hypergraph with hyperedges $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$, then we can say if $|\mathbf{e}_{i}|<|\mathbf{e}_{j}|$ and $\mathbf{e}_{i}\subset\mathbf{e}_{j}$, then $w_{\mathbf{e}_{j}}\leq w_{\mathbf{e}_{i}}$. This means that higher cardinality hyperedges will be assigned a weight that is either the same or less than the weight of some lower cardinality hyperedge subset. Therefore, if the selection of hyperedges prioritizes choosing those with the largest weights first, then the closed form solution implies that lower cardinality hyperedges could be assigned greater importance. This is an assumption that may not hold in all cases, and establishes more priors beyond that of smoothness.
@@
The second shortcoming of HGSI is in the method of hyperedge possibility reduction. The use of K-NN provides a notion of local smoothness on the hypergraph by focusing on one node at a time and its nearest neighbors with no other outside information. While local smoothness is certainly a related concept, the use of K-NN fails to capture any global smoothness which is of greater interest when establishing a smoothness prior. This reduction method exaggerates the overlap between a hypergraph’s true hyperedges and those produced by K-NN. HGSI ends up discarding valid hyperedge possibilities under the local smoothness assumption, as we will demonstrate in Sec. V.
IV Hypergraph Structure Learning with Smoothness
@@
We now formulate our own method to recover the hypergraph structure to overcome the shortcomings of previous works. Our method, which we coin as Hypergraph Structure Learning with Smoothness (HSLS), centers around learning the weight vector $\mathbf{w}\in\mathbb{R}^{D}$ for an initial $D$ hyperedge possibilities. Using the work of Tang et al. in (9) as reference, we define our total variation as $\Omega(\mathbf{w})=\mathbf{z}^{\top}\mathbf{w}$. In the same experimental context, the elements of $\mathbf{z}$ will vary depending on the smoothness term selected, but it always contains the value of the time-series difference measures. So, for the Sum-Square total variation, $\mathbf{z}$ holds the sum of the squared differences, while for Max-Absolute, it holds the maximum of the absolute differences, and so on for the other terms. The format of $\Omega(\mathbf{w})$ provides the ability to substitute in any of the selected smoothness terms creating a robust optimization. Consequently, we can evaluate in the same setting which terms perform better. Furthermore, we reiterate that we invoke a smoothness prior for the signals on the hypergraph structure which $\Omega(\mathbf{w})$ enforces.
@@
Revisiting the general form of (1), the use of $\theta(\mathbf{w})$ includes other terms that impose further structure on the hypergraph. We will be using a degree positivity and sparsity term in our $\theta(\mathbf{w})$. After substituting the terms into (1), the final formulation of our optimization problem is
@@
where $\log(\cdot)$ is the natural logarithm taken element-wise and $||\cdot||_{2}^{2}$ is the squared l2 norm of a vector. The $\alpha$ and $\beta$ are positive scalar parameters used to control the influence of their corresponding function. The matrix $\mathbf{S}\in\mathbb{R}^{N\times D}$ performs a linear transformation on $\mathbf{w}$ to convert $\mathbf{Sw}=\mathbf{d}$. Here, $\mathbf{d}\in\mathbb{R}^{N}$ is the degree vector containing the degree of each node which is the sum off all the weights corresponding to hyperedges that contain $v_{i}$. Hence, the negative log term acts on degree positivity by allowing nodes to still have a degree close to zero, but remain positive. Also, since it is acting on $\mathbf{d}$, the $\alpha$ parameter can be used to control the magnitude of the node degrees. The squared l2 norm is the sparsity term which controls the spread of the weights across the hyperedges. Similar to $\alpha$, the $\beta$ parameter is used to control the level of influence of sparsity on the hypergraph structure.
Here, we note the key differences between (9) and (11). First, as previously mentioned, our smoothness term is no longer fixed, but allowed to be substituted with the selected definitions. Second, the degree positivity term is now based on $\mathbf{d}$ instead of strictly $\mathbf{w}$. The inclusion of $\mathbf{d}$ is used in the tensor based hypergraph learning work of Pena-Pena et al. and the graph learning work of Kalofolias. The use of the degree vector is critical to avoid the closed form solution of (10), which in turn overcomes the issue of hyperedges being assigned unbalanced importance. Third, we change the sparsity term to be the squared l2 norm which makes the optimization more sensitive to smaller perturbations in $\mathbf{w}$. Finally, we include the parameters $\alpha$ and $\beta$ for greater control over the influence of each structural parameter.
@@
The only new term we have introduced is the indicator function $\mathbbm{1}\{\cdot\}$ where we define it as
$\mathbbm{1}\{w_{i}\geq 0\}=\left\{\begin{aligned} &1\quad&,w_{i}\geq 0\\ &+\infty\quad&,w_{i}<0\end{aligned}\right.$ (16)
@@
From Komodakis and Pesquet, we have chosen to use the Forward-Backward-Forward (FBF) based primal-dual algorithm to perform the minimization. The FBF algorithm requires the use of the gradient of $h(\cdot)$, the proximal operator of $f(\cdot)$, and the proximal operator of the conjugate of $g(\cdot)$ denoted as $g^{*}(\cdot)$. These operators are defined as
@@
Now, in reference to the size of the hyperedge search space, the use of $D$ hyperedge possibilities for every potential hyperedge permutation is incredibly costly as $N$ and the number of elements of $\mathcal{K}$ increases. The HGSI method attempted to counteract this by reducing the hyperedge selection set using each node’s $k_{l}-1$ nearest neighbors. This approach is not very effective as many potential valid hyperedges are removed from consideration. We demonstrate these concepts in Sec. V where we find that, after a certain number of nearest neighbors, all hyperedges of a ground truth hypergraph are in the valid hyperedge set. So, we adopt an approach where we increase the number of nearest neighbors for each node across each cardinality $k_{l}$ resulting in $\mathbf{w}\in\mathbb{R}^{\hat{D}}$ for $\hat{D}$ number of reduced hyperedge possibilities. This method of hyperedge reduction is detailed in Algorithm 1 which operates with $\mathbf{X}$ being the hypergraph signals across a set of nodes $V(\mathcal{H})$.
Algorithm 1 Method of Hyperedge Reduction for HSLS
1:  Input: $\mathbf{X},\hat{\mathcal{K}},V(\mathcal{H}),$
@@
3:  for $\hat{k}_{l}\in\hat{\mathcal{K}}$ do
4:     for $\mathbf{v}_{i}\in V(\mathcal{H})$ do
5:        $\mathcal{J}=KNN(\mathbf{X},v_{i},\hat{k}_{l})$
6:        Create hyperedge sets $\mathbf{e}_{d}$ from $v_{i}$ combined with every unique permutation of $\hat{k}_{l}$ nodes from $\mathcal{J}$
7:        Store each new $\mathbf{e}_{d}$ into reduced hyperedge set $\hat{E}(\mathcal{H})$
8:     end for
@@
From an initial set of cardinalities $\mathcal{K}$, the elements $\hat{k}_{l}\in\hat{\mathcal{K}}$ are such $\hat{k}_{l}\geq k_{l}-1$ and represent the number of neighbors per cardinality in $\mathcal{K}$ to keep for each node. If $\hat{k}_{l}=k_{l}-1$, then this degrades into the exact same process as from HGSI.
We define function $KNN(\cdot)$ to produce a set of $\hat{k}_{l}$ nodes that are the nearest neighbors of $v_{i}$ based on Euclidean distances of the time-series in $\mathbf{X}$. The output $\hat{E}(\mathcal{H})$ indicates the set of all hyperedges recovered using nearest neighbors, and $\hat{D}=|\hat{E}(\mathcal{H})|$ is the number of hyperedges in the set. In our work, we have a guided method of selecting values in $\hat{\mathcal{K}}$; however, in applications without underlying hypergraph information, this selection must be based on reasonable speculation. Algorithm 1 avoids the issues of the hyperedge possibility reduction assumption of HGSI while simultaneously increasing our algorithm’s scalability.
Algorithm 2 Hypergraph Structure Learning with Smoothness (HSLS): FBF Primal-Dual Algorithm for solving
@@
We can now define our final model for HSLS as shown in Algorithm 2. Most of the inputs are the same as those included in (11) where $\mathbf{w}$ is the primal variable and $\mathbf{d}$ is the dual variable. With the inclusion of Algorithm 1, we updated the appropriate dimensions with $\hat{D}$. The superscripts indicate which iteration of the vectors are being used as they are updated on each pass. The parameter, $\gamma_{i}$, is effectively the learning rate, but in every forward pass, it is updated according to a sequence determined by the Lipschitz constant, $\zeta$. See Appendix C for details on the derivation of this sequence. The $\eta$ is the threshold for the stopping condition if the max iteration $i_{max}$ is not reached first. Following the guidelines of Komodakis and Pesquet, HSLS is guaranteed to converge.
@@ V Experiments
We now evaluate the performance of our model in experimental settings [^1] All source code used for our experiments is publicly available at https://github.com/Ben-Brown-Code/Scalable-Hypergraph-Structure-Learning-with-Diverse-Smoothness-Priors. First, we test on co-authorship networks with both uniform and non-uniform structures. In this scenario, the underlying hypergraph structure is known and will serve as our ground truth. The purpose of this experiment is to evaluate the accuracy with which our optimization combined with various total variations can recover a hypergraph topology. The results are a report of F1-Score, Precision, and Recall between the learned hypergraph and the ground truth. Furthermore, this experiment serves to demonstrate the shortcomings of HGSI using K-NN as a method of hyperedge possibility reduction, as mentioned in Sec. III, and how that method can be improved upon. We then test on a set of lung cancer mortality data across the U.S. This experiment is designed to evaluate the effectiveness of each total variation in producing a smooth hypergraph structure. This data has no ground truth, but a comparison to a baseline hypergraph is used to highlight the emphasis our algorithm puts on selecting hyperedges that create a globally smooth topology. Finally, we use a large time-series dataset of voltage readings to test the speed and scalability of our proposed method.

### v2

@@ III Hypergraph Structure Inference Under Smoothness Prior (HGSI)
calculated element-wise on $\mathbf{z}$ where $\mathbf{w}^{\star}\in(0,1]^{\bar{D}}$. The recovery of $\mathbf{w}^{\star}$ is the equivalent of learning the hypergraph structure since it contains the weights for each potential hyperedge, indicating a level of confidence that each hyperedge is in the hypergraph.
@@
We note that the length of $\mathbf{w}^{\star}$ is not $D$, but rather $\bar{D}$. It is recognized that computing $\mathbf{w}^{\star}$ is time inefficient and, as $N$ and $M$ increase, often completely infeasible due to the vast number of hyperedge combinations. So, to limit the number of hyperedge possibilities, HGSI employs the use of a K-NN based method for possibility reduction. Specifically, given the set of cardinalities $\mathcal{K}$, determine the $k_{l}-1$ nearest neighbors of each node based on time-series observations. Then, let the combinations of the original nodes and their nearest neighbors be the new set of possible hyperedges where $\bar{D}=NL$ is the upper bound of hyperedge possibilities in the set reduced by K-NN for $L=|\mathcal{K}|$. This is considered an upper bound since the use of K-NN could produce repeat hyperedges which would eventually be aggregated into one copy.
@@
HGSI has advantages when it comes to scalability through the use of a closed form solution and the hyperedge reduction technique. However, both of these factors also cause certain shortcomings as they relate to assumptions on hypergraph structure. The first is that the format of (11) improperly constrains the set of learned hyperedges in a non-uniform hypergraph structure as a factor of $\mathbf{z}$. If there is some non-uniform hypergraph with hyperedges $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$, then we can say if $|\mathbf{e}_{i}|<|\mathbf{e}_{j}|$ and $\mathbf{e}_{i}\subset\mathbf{e}_{j}$, then $w_{\mathbf{e}_{j}}\leq w_{\mathbf{e}_{i}}$. This means that higher cardinality hyperedges will be assigned a weight that is either the same or less than the weight of some lower cardinality hyperedge subset. Therefore, if the selection of hyperedges prioritizes choosing those with the largest weights first, then the closed form solution implies that lower cardinality hyperedges could be assigned greater importance. This is an assumption that may not hold in all cases, and establishes more priors beyond that of smoothness.
@@
The second shortcoming of HGSI is in the method of hyperedge possibility reduction. The use of K-NN provides a notion of local smoothness on the hypergraph by focusing on one node at a time and its nearest neighbors with no other outside information. While local smoothness is certainly a related concept, the use of K-NN fails to capture any global smoothness which is of greater interest when establishing a smoothness prior. This reduction method exaggerates the overlap between the set of true hyperedges and those produced by K-NN. HGSI ends up discarding valid hyperedge possibilities under the local smoothness assumption, as we will demonstrate in Sec. V.
IV Hypergraph Structure Learning with Smoothness (HSLS)
We now formulate our own method to recover the hypergraph structure to overcome the shortcomings of previous works. Our method, which we coin as Hypergraph Structure Learning with Smoothness (HSLS), centers around learning the weight vector $\mathbf{w}\in\mathbb{R}^{D}$ for an initial $D$ hyperedge possibilities. It is important to emphasize that the optimization procedure that will be implemented is from an existing source that is fit to our model. Although the problem will be shown as being solved using a standard convex optimization algorithm, the cost function is newly implemented. This cost function encodes domain-specific structure that has not been captured in prior formulations, and it is one source of the methodological novelty. Furthermore, while the employed convex optimization algorithm is pre-existing, since the cost function is a new interpretation, this constitutes a unique optimization setup.
@@ IV Hypergraph Structure Learning with Smoothness (HSLS)
Using the work of Tang et al. in (10) as reference, we define our total variation as $\Omega(\mathbf{w})=\mathbf{z}^{\top}\mathbf{w}$. In the same experimental context, the elements of $\mathbf{z}$ will vary depending on the smoothness term selected, but it always contains the value of the time-series difference measures. So, for the Sum-Square total variation, $\mathbf{z}$ holds the sum of the squared differences, while for Max-Absolute, it holds the maximum of the absolute differences, and so on for the other terms. The format of $\Omega(\mathbf{w})$ provides the ability to substitute in any of the selected smoothness terms creating a robust optimization. Consequently, we can evaluate in the same setting which terms perform better. Since all total variation terms can fit this convex format, the optimization we pose will also maintain convexity. Furthermore, we reiterate that we invoke a smoothness prior for the signals on the hypergraph structure which $\Omega(\mathbf{w})$ enforces.
@@
Revisiting the general form of (1), the use of $\theta(\mathbf{w})$ includes other terms that impose further structure on the hypergraph. We will be using a degree positivity and sparsity term in our $\theta(\mathbf{w})$. After substituting the terms into (1), the final formulation of our convex cost function is
@@
where $\log(\cdot)$ is the natural logarithm taken element-wise and $||\cdot||_{2}^{2}$ is the squared $\ell_{2}$ norm of a vector. The $\alpha$ and $\beta$ are positive scalar parameters used to control the influence of their corresponding function. The matrix $\mathbf{S}\in\mathbb{R}^{N\times D}$ performs a linear transformation on $\mathbf{w}$ to convert $\mathbf{Sw}=\mathbf{d}$. Here, $\mathbf{d}\in\mathbb{R}^{N}$ is the degree vector containing the degree of each node which is the sum off all the weights corresponding to hyperedges that contain $v_{i}$. Hence, the negative log term acts on degree positivity by allowing nodes to still have a degree close to zero, but remain positive. Also, since it is acting on $\mathbf{d}$, the $\alpha$ parameter can be used to control the magnitude of the node degrees. The squared $\ell_{2}$ norm is the sparsity term which controls the spread of the weights across the hyperedges. Similar to $\alpha$, the $\beta$ parameter is used to control the level of influence of sparsity on the hypergraph structure.
Here, we note the key differences between (10) and (12). First, as previously mentioned, our smoothness term is no longer fixed, but allowed to be substituted with the selected total variation definitions. Second, the degree positivity term is now based on $\mathbf{d}$ instead of strictly $\mathbf{w}$. The inclusion of $\mathbf{d}$ is used in the tensor based hypergraph learning work of Pena-Pena et al. [26] and the graph learning work of Kalofolias [16]. The use of the degree vector is critical to avoid the closed form solution of (11), which in turn overcomes the issue of hyperedges being assigned unbalanced importance. Third, we change the sparsity term to be the squared $\ell_{2}$ norm which makes the optimization more sensitive to smaller perturbations in $\mathbf{w}$. Finally, we include the parameters $\alpha$ and $\beta$ for greater control over the influence of each structural parameter.
@@
The only new term we have introduced is the indicator function $\mathbbm{1}\{\cdot\}$ defined as
$\mathbbm{1}\{w_{\mathbf{e}_{d}}\geq 0\}=\left\{\begin{aligned} &1\quad&,w_{\mathbf{e}_{d}}\geq 0\\ &+\infty\quad&,w_{\mathbf{e}_{d}}<0\end{aligned}\right.$ (17)
@@
From Komodakis and Pesquet [47], we have chosen to use the convex Forward-Backward-Forward (FBF) based primal-dual algorithm to perform the minimization. The FBF algorithm requires the use of the gradient of $h(\cdot)$, the proximal operator of $f(\cdot)$, and the proximal operator of the conjugate of $g(\cdot)$ denoted as $g^{*}(\cdot)$. These operators are defined as
@@
Now, in reference to the size of the hyperedge search space, the use of $D$ hyperedge possibilities, which is every potential hyperedge combination, is incredibly costly as $N$ and $|\mathcal{K}|$ increase. HGSI attempts to counteract this by reducing the hyperedge selection set using each node’s $k_{l}-1$ nearest neighbors. This approach is not very effective as many potential valid hyperedges are removed from consideration. We demonstrate these concepts in Sec. V where we find that, after a certain number of nearest neighbors, all hyperedges of a ground truth hypergraph can be included in the valid hyperedge set.
Algorithm 1 Method of Hyperedge Reduction for HSLS.
1:  Input: $\mathbf{X},{\mathcal{K}},{\mathcal{R}},V(\mathcal{H}),$
@@
3:  for $l=1:L$ do
4:     for $i=1:N$ do
5:        $\mathcal{J}_{v_{i}}=KNN(\mathbf{X},v_{i}\in V(\mathcal{H}),r_{l}\in{\mathcal{R}})$ set of $r_{l}$ nodes that are nearest neighbors of $v_{i}$ in $\mathbf{X}$, $v_{i}$ excluded.
6:        $\mathcal{C}_{v_{i}}=\{\{v_{i},v_{j_{1}},v_{j_{2}},\dots,v_{j_{k_{l}-1}}\}:v_{j_{p}},v_{j_{q}}\in\mathcal{J}_{v_{i}},v_{j_{p}}\neq v_{j_{q}},k_{l}\in\mathcal{K}\}$ all unique combinations of $v_{i}$ with $k_{l}-1$ nodes from $\mathcal{J}_{v_{i}}$.
7:     end for
8:     $\hat{E}_{k_{l}}(\mathcal{H})=\bigcup_{i=1}^{N}\mathcal{C}_{v_{i}}$ reduced set of hyperedges for cardinality $k_{l}$.
@@
10:  $\hat{E}(\mathcal{H})=\bigcup_{l=1}^{L}\hat{E}_{k_{l}}(\mathcal{H})$ reduced set of hyperedges for all cardinalities.
To this end, we adopt an approach where a variable number of nearest neighbors can be selected per cardinality in $\mathcal{K}$. This method of hyperedge reduction can be found in Algorithm 1 and is detailed here. We introduce =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{L}\}$ such that $r_{l}\geq k_{l}-1$ is a lower bound, and $|\mathcal{R}|=|\mathcal{K}|$. The elements of set $\mathcal{R}$ are considered the number of nearest neighbors to be taken per node with reference to $\mathcal{K}$. Clarifying this point, $\mathcal{K}$ is the set of cardinalities in a hypergraph. Each $r_{l}$ corresponds to one of these cardinalities $k_{l}$ specifying the number of nearest neighbors in $\mathbf{X}$ to be taken. The $r_{l}$ nearest neighbors of node $v_{i}$ get stored in $\mathcal{J}_{v_{i}}$. We define function $KNN(\cdot)$ to produce the $r_{l}$ nodes that are the nearest neighbors of $v_{i}$ based on Euclidean distances of the time-series in $\mathbf{X}$. We then take every unique combination of $v_{i}$ with $k_{l}-1$ nodes from $\mathcal{J}_{v_{i}}$ and store as set $\mathcal{C}_{v_{i}}$. The union of all $\mathcal{C}_{v_{i}}$ produces $\hat{E}_{k_{l}}(\mathcal{H})$ the set of hyperedges for one cardinality, and the union of $\hat{E}_{k_{l}}(\mathcal{H})$ across all cardinalities is $\hat{E}(\mathcal{H})$, the final reduced hyperedge set. This hyperedge set has an upper bound size
$\hat{D}=\sum_{l=1}^{L}\binom{r_{l}}{k_{l}-1},$ (21)
which is considered an upper bound since the use of K-NN could produce repeat hyperedges which are aggregated into one copy, similar to the discussion of HGSI.
To make this more concrete, consider the following example. Let $\mathcal{K}=\{3,4\}$ and =$\mathcal{R}=\{15,10\}$ for $N=10$. Then, Algorithm 1 would start with $r_{1}=15$ which corresponds to cardinality $k_{1}=3$. Then, considering only $v_{1}$ currently, its $r_{1}=15$ nearest neighbors from $\mathbf{X}$ (excluding itself) are found and labeled set $\mathcal{J}_{v_{1}}$. From the $r_{1}=15$ nodes in $\mathcal{J}_{v_{1}}$, $v_{1}$ is paired with every unique combination of $k_{1}-1=2$ of these nearest neighbor nodes. This produces $\binom{r_{1}}{k_{1}}=105$ combinations of order $k_{1}=3$ hyperedges between $v_{1}$ and its set of $r_{1}=15$ nearest neighbors from $\mathcal{J}_{v_{1}}$, and this is stored in $\mathcal{C}_{v_{1}}$. This process is repeated for every other node to get $\hat{E}_{k_{1}}(\mathcal{H})=\bigcup_{i=1}^{10}\mathcal{C}_{v_{i}}$, and then again for every other cardinality in $\mathcal{K}$ for final output $\hat{E}(\mathcal{H})=\bigcup_{l=1}^{2}\hat{E}_{k_{l}}(\mathcal{H})$.
Algorithm 2 FBF Primal-Dual Algorithm for solving
@@
We can now define the minimization solver for (12) as shown in Algorithm 2. Most of the inputs are the same as those included in (12) where $\mathbf{w}$ is the primal variable and $\mathbf{d}$ is the dual variable. With the inclusion of Algorithm 1 as a pre-processing step for Algorithm 2, we updated the appropriate dimensions with $\hat{D}$. The superscripts indicate which iteration of the vectors are being used as they are updated on each pass. The parameter, $\gamma_{i}$, is effectively the learning rate, but in every forward pass, it is updated according to a sequence determined by the Lipschitz constant, $\zeta$. See Appendix C for details on the derivation of this sequence. The $\eta$ is the threshold for the stopping condition if the max iteration $i_{max}$ is not reached first. The convergence rate is $\mathcal{O}(1/(i+1))$ for iteration $i$, and the overall complexity is $\mathcal{O}(N\hat{D})$. The proximal operator of function $f$ has complexity $\mathcal{O}(\hat{D})$ and the proximal operator of function $g^{*}$ has complexity $\mathcal{O}(N)$. Following the guidelines of Komodakis and Pesquet [47], HSLS is guaranteed to converge to the approximate global minimum.
@@ V Experiments
We now evaluate the performance of our model in experimental settings [^1] All source code used for our experiments is publicly available at https://github.com/Ben-Brown-Code/Scalable-Hypergraph-Structure-Learning-with-Diverse-Smoothness-Priors. First, we test on co-authorship networks with both uniform and non-uniform structures. In this scenario, the underlying hypergraph structure is known and will serve as our ground truth. The purpose of this experiment is to evaluate the accuracy with which our optimization combined with various total variations can recover a hypergraph topology. The results are a report of F1-Score, Precision, and Recall between the learned hypergraph and the ground truth. Furthermore, this experiment serves to demonstrate the shortcomings of HGSI using K-NN as a method of hyperedge possibility reduction, as mentioned in Sec. III, and how that method can be improved upon. We then test on a set of lung cancer mortality data across the U.S. This experiment is designed to evaluate the effectiveness of each total variation in producing a smooth hypergraph structure. This data has no ground truth, but a comparison to a baseline hypergraph is used to highlight the emphasis our algorithm puts on selecting hyperedges that create a globally smooth topology. Finally, we apply HSLS to a hyperspectral image classification task to improve the results of a baseline method and further the total variation analysis.